The Faraway Moving Company is involved in a major plant expansion that involves the expenditure of ​$207 million in the coming year. The firm plans on financing the expansion through the retention of ​$139 million in firm earnings and by borrowing the remaining ​$68 million. In return for helping sell the ​$68 million in new​ debt, the​ firm's investment banker charges a fee of 250 basis points​ (where one basis point is 0.01​ percent). If Faraway decides to adjust for these flotation costs by adding them to the initial​ outlay, what will the initial outlay for the project​ be?

The flotation cost adjusted initial outlay is $
Answer #1

Flotation cost = Net Borrowed amount * Flotation fee/(1-Flotation fee)

= 68000000*2.5%/(1-2.5%)

= 1743589.74

The flotation cost adjusted initial outlay = Total expense+flotation cost

= 207000000+1743589.74 =

208743589.74